hey can you help me... the name of the chapter is: solving problems involving quadratic equations.
question :
the length of a rectangle is 5m more than its width. its area is 234m squared ,find the dimension of the rectangle

Answer by checkley79(3341)

AREA=LW
L=W+5
234=(W+5)W
234=W^2+5W
W^2+5W-234=0
(W+18)(W-13)=0
W-13=0
W=13 ANS. FOR THE WIDTH.
L=13+5=18 ANS. FOR THE LENGTH.
PROOF:
234=13*18
234=234